Grace is twice as old as Tim. Ten years ago, she was 4 times as old as Tim. How old is each now?
Answer by stanbon(75887)

Grace is twice as old as Tim. Ten years ago, she was 4 times as old as Tim. How old is each now?
-------------
Equations:
g = 2t
(g-10)=4(t-10)
---------------------
Substitute for "g" and solve for "t":
(2t-10) = 4t-40
2t = 30
t = 15 (Tim's age now)
---
Since g = 2t, g = 30 (Grace's age now)
==========================================
Cheers,
Stan H.